OpenStudy (anonymous):

Did I get this right? How many real-number solutions does the equation have? 7x^2 + 8x + 5 = 0

OpenStudy (anonymous):

one solution two solutions** no solutions infinitely many solutions

jimthompson5910 (jim_thompson5910):

what did you get?

jimthompson5910 (jim_thompson5910):

oh nvm, one sec while I check

jimthompson5910 (jim_thompson5910):

no that's not correct

jimthompson5910 (jim_thompson5910):

the discriminant is D = b^2 - 4ac

jimthompson5910 (jim_thompson5910):

This determines what type and how many solutions you'll have

jimthompson5910 (jim_thompson5910):

D = b^2 - 4ac D = 8^2 - 4(7)(5) D = 64 - 140 D = -76 because the discriminant is negative, you'll have no real solutions

jimthompson5910 (jim_thompson5910):

The solutions instead will be complex or imaginary (if you haven't heard of either term, then don't worry)

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

@jim_thompson5910 Could you help with one more?

OpenStudy (anonymous):

A model rocket is launched from a roof into a large field. The path of the rocket can be modeled by the equation y = 0.04x+ 8.3x + 4.3, where x is the horizontal distance, in meters, from the starting point on the roof and y is the height, in meters, of the rocket above the ground. How far horizontally from its starting point will the rocket land? 208.02 m 416.03 m 0.52 m 208.19 m

jimthompson5910 (jim_thompson5910):

is there an x^2 in there?

OpenStudy (anonymous):

y = 0.04x^2+ 8.3x + 4.3 Yes sorry

jimthompson5910 (jim_thompson5910):

and is the leading coefficient negative?

OpenStudy (anonymous):

Yeah, woops

jimthompson5910 (jim_thompson5910):

ok so you you'll need to find the roots of y = -0.04x^2 + 8.3x + 4.3

jimthompson5910 (jim_thompson5910):

the best way to do so is to use the quadratic formula

jimthompson5910 (jim_thompson5910):

have you heard of that formula before?

OpenStudy (anonymous):

So that's ax+b+c ? Not sure if that's right.

jimthompson5910 (jim_thompson5910):

The quadratic formula is \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

you would plug a = -0.04, b = 8.3 and c = 4.3 in to get \[\Large x = \frac{-(8.3)\pm\sqrt{(8.3)^2-4(-0.04)(4.3)}}{2(-0.04)}\]

OpenStudy (anonymous):

I'm still confused..

jimthompson5910 (jim_thompson5910):

From this point onward, it's just a matter of using a calculator to evaluate and simplify \[\Large x = \frac{-8.3\pm\sqrt{68.89-(-0.688)}}{-0.08}\] \[\Large x = \frac{-8.3\pm\sqrt{68.89+0.688}}{-0.08}\] \[\Large x = \frac{-8.3\pm\sqrt{69.578}}{-0.08}\] \[\Large x = \frac{-8.3+\sqrt{69.578}}{-0.08} \ \text{or} \ x = \frac{-8.3-\sqrt{69.578}}{-0.08}\] \[\Large x = \frac{-8.3+\sqrt{69.578}}{-0.08} \ \text{or} \ x = \frac{-8.3-\sqrt{69.578}}{-0.08}\] \[\Large x = -.5167852195 \ \text{or} \ x = 208.0167852\] \[\Large x \approx -0.516785 \ \text{or} \ x \approx 208.016785\]

jimthompson5910 (jim_thompson5910):

it's a bit of a mess, but in the end, the two roots are approximately x = -0.516785 or x = 208.016785

OpenStudy (anonymous):

Is it a?

jimthompson5910 (jim_thompson5910):

ignore the negative root and focus only on the positive root

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

going from x = 0 to x = 208.016785 is roughly 208.02 ft

OpenStudy (anonymous):

Alright. Thanks again!

jimthompson5910 (jim_thompson5910):

sure thing